OpenStudy (anonymous):

A wheel, originally rotating at 126 rad/s undergoes a constant angular deceleration of 5.00 rad/s2. What is its angular speed after it has turned through an angle of 628 radians?

OpenStudy (anonymous):

Just like the linear kinematic equations \[ \omega^2 = \omega_o^2 + 2 \alpha \theta\] \[ \omega = \ \textrm{final angular velocity} \\ \omega_o = \ \textrm{initial angular velocity} \\ \alpha = \ \textrm{angular acceleration} \\ \theta = \ \textrm{angular displacement} \]

OpenStudy (anonymous):

Yep I use that formula but I didnot get right answer

OpenStudy (anonymous):

;/ What answer did you get?

OpenStudy (anonymous):

\[ \alpha\] is negative ^_^

OpenStudy (anonymous):

Why is that negative?

OpenStudy (anonymous):

Acceleration is proportional to displacement but acts in opposite direction to the displacement Displacement is modeled by the sin graph, differentiate that to get velocity = cos graph, differentiate that to get acceleration = -sin graph

OpenStudy (anonymous):

so they are always negative in any case right?

OpenStudy (anonymous):

It's negative in this case because it's a "deceleration" - a negative acceleration. The other way of saying it in this problem would be " an acceleration of -5.00 rad/s"

OpenStudy (anonymous):

~~rad/s^2

OpenStudy (anonymous):

The graph of the acceleration is always phase shifted by pi in relation to the angular displacement. The magnitude of the acceleration is dependent on the problem.

OpenStudy (anonymous):

thank so much I got it